# UI-AGILE: Advancing GUI Agents with Effective Reinforcement Learning and Precise Inference-Time Grounding

### v2

@@ 4.3 Agent Capability Evaluation
Table 4: Action type accuracy, grounding accuracy and success rate on AndroidControl-Low and AndroidControl-High. Results marked in bold represent the best performance, and underlined results indicate the second-best performance.
@@
Following OS-Atlas, we use 7,708 examples for a fair comparison, while some works (e.g., Aguvis (Xu et al. 2024) and UGround (Gou et al. 2025)) randomly sample 500 action steps for testing. The evaluation is conducted under two distinct settings. In the AndroidControl-Low setting, the agent receives a specific, low-level instruction for each step. In contrast, the AndroidControl-High setting provides the agent with a high-level goal, requiring it to infer the correct action for the current step based on the conversation history and previous actions, thereby testing its multi-step reasoning capability.
As presented in Tab. 4, our UI-AGILE-7B model achieves the best performance (SR: 77.6 and 60.6) compared to other RFT models including UI-R1-E (SR: 71.37 and 35.88), GUI-R1-3B (SR: 64.41 and 46.55) and GUI-R1-7B (SR: 66.52 and 51.67).
Notably, our smaller UI-AGILE-3B model with SR of 77.4 and 56.9 also surpasses 7B baselines like GUI-R1-7B (SR: 66.5 and 51.7). For Type, a similar trend can be observed.
@@ 4.4 Ablation Study
To further verify the contribution of each of our training techniques (“Simple Thinking” reward, continuous grounding reward, and cropping-based resampling), we conduct an ablation study of UI-AGILE-3B and the results are shown in Fig. 3 and Fig. 4. We can observe that:
  • Applying continuous grounding reward and cropping-based resampling improve the performance by approximately 10% and 12.4% on ScreenSpot-Pro, respectively. The former incentivizes more precise localization to the target’s center and the later helps avoid ineffective training with zero reward. They also slightly improve grounding accuracy on ScreenSpot-v2 where the performance of the base model is already high and it is difficult to achieve significant gains.
  • Removing “Simple Thinking” during training (i.e., “No Thinking”) leads to higher grounding accuracy on ScreenSpot-Pro and ScreenSpot-v2 (about 0.4% and 0.7% gains). However, Fig. 4 demonstrates that integrating “Simple Thinking” enhances the agent’s decision-making with planning, resulting in a noticeable improvement in SR on the AndroidControl benchmark, with SR increases of 15.5% and 3.4% in Low and High settings, respectively.
Figure 3: Ablation study on grounding benchmarks.
Figure 4: Ablation study on AndroidControl benchmark.
4.5 Analysis of Attempts Per Step
[…]
@@ 4.6 Analysis of Inference Time
As a baseline, the standard grounding approach applied to UI-AGILE-7B without our method completes the benchmark in 30 minutes. When applying our method, the decomposed grounding stage takes 35 minutes. The subsequent VLM-based selection stage requires additional 4 minutes. The modest increase in overhead is a practical trade-off for the substantial gain of grounding accuracy brought by our method.
@@ 5 Conclusions
In this paper, we introduce UI-AGILE, a comprehensive framework designed to enhance GUI agents’ training and inference. We tackle the practical challenges of the reasoning-grounding dilemma, ineffective reward, and visual noise. During training, our solution integrates three key innovations: a “Simple Thinking” reward to foster efficient yet effective reasoning, a continuous grounding reward that incentivizes high-precision localization, and a cropping-based resampling strategy to overcome the sparse reward problem. For inference, we introduce decomposed grounding with selection, a novel method that reduces visual noise and dramatically improves grounding accuracy on high-resolution screens while the inference cost is only slightly increased. Experimental results demonstrate the effectiveness of our proposed techniques on enhancing GUI agents’ grounding capability and general capability.
Limitations and future work.
The VLM used in the selection stage of our decomposed grounding with selection method is a general-purpose, pre-trained model. A promising future direction would be to fine-tune this adjudicator model on a curated dataset of candidate UI elements, enhancing its selection accuracy to achieve further gains in overall grounding performance.

### v3

@@ 4.3 Agent Capability Evaluation
Table 3: Type, GR and SR on AndroidControl-Low and AndroidControl-High. Results marked in bold represent the best performance, and underlined results indicate the second-best performance.
@@
Following OS-Atlas, we use 7,708 examples for a fair comparison, while some works (e.g., Aguvis (Xu et al. 2024) and UGround (Gou et al. 2025)) randomly sample 500 action steps for testing. The evaluation is conducted under two distinct settings. In AndroidControl-Low, the agent receives a specific, low-level instruction for each step. In contrast, AndroidControl-High provides the agent with a high-level goal, requiring it to infer the correct action for the current step based on the conversation history and previous actions, thereby testing its multi-step reasoning capability.
As depicted in Tab. 3, UI-AGILE-7B achieves the best performance (SR: 77.6 and 60.6) compared to other RFT models, including UI-R1-E (SR: 71.37 and 35.88), GUI-R1-3B (SR: 64.41 and 46.55) and GUI-R1-7B (SR: 66.52 and 51.67). Notably, the smaller UI-AGILE-3B (SR: 77.4 and 56.9) also surpasses 7B baselines like GUI-R1-7B (SR: 66.5 and 51.7). For Type, a similar trend can be observed.
@@ 4.4 Ablation Study
Figure 3: Ablation study.
To verify the contribution of each training technique, we conduct an ablation study of UI-AGILE-3B and the results are shown in Fig. 3(a) and Fig. 3(b). Due to page limit, the performance on AndroidControl w.r.t. Type and GR are provided in Appendix C. We can observe that:
  • Applying continuous grounding reward and cropping-based resampling improves the performance by 10% and 12.4% on ScreenSpot-Pro, respectively. The former incentivizes more precise localization to the target’s center and the later helps avoid ineffective training with zero reward. They also slightly improve grounding accuracy on ScreenSpot-v2 where the performance of base model is already high and it is difficult to achieve significant gains.
  • Removing “Simple Thinking” during training (i.e., “No Thinking”) leads to higher grounding accuracy on ScreenSpot-Pro and ScreenSpot-v2 (0.4% and 0.7% gains). However, Fig. 3(b) shows that integrating “Simple Thinking” enhances the agent’s decision-making with planning, resulting in a noticeable improvement in SR on AndroidControl, with SR increases of 15.5% and 3.4% in Low and High settings, respectively.
@@ 4.5 Analysis of Inference Time
As a baseline, the standard grounding approach applied to UI-AGILE-7B completes the benchmark in 30 minutes. When applying our method, the decomposed grounding stage takes 35 minutes. The subsequent VLM-based selection stage requires additional 4 minutes. The modest increase in overhead is a practical trade-off for the substantial gain of grounding accuracy brought by our method.
4.6 Analysis of Attempts Per Step
Due to space limit, we include the analysis of attempts per training step in Appendix D.
@@ 5 Conclusions
In this paper, we introduce UI-AGILE, a comprehensive framework designed to enhance GUI agents’ training and inference. It tackles the practical challenges of the reasoning-grounding dilemma, ineffective reward, and visual noise. Experimental results demonstrate the effectiveness of our proposed techniques on enhancing GUI agents’ grounding capability and general capability.
Despite the promising performance of UI-AGILE, the VLM used in its decomposed grounding with selection method is a general-purpose, pre-trained model that is not originally tuned for the selection task. A promising future direction would be to fine-tune this adjudicator model on a curated dataset of candidate UI elements, enhancing its selection accuracy to achieve further gains in overall grounding performance.
Appendix A Environment and Implementation Details
In the following, we provide the details of experiment environment and our implementation.
Experiment Environment.
We use a machine with two Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz, 128GB main memory and eight NVIDIA A800 GPU for experiments.
Training Details.
We use the trl framework to implement the cropping-based resampling strategy and reward functions. The sampling process is attempted 4 times at most and is bypassed entirely if the bbox’s dimensions exceed the target crop size. Following prior works, we use Qwen2.5-VL-3B and Qwen2.5-VL-7B as base models.
Inference Details.
For decomposed grounding with selection, the input image is divided into four sub-images scaling to 60% of the original dimensions, with adjacent sub-images overlapping by 10% of the original image’s width and height. In the element image extraction stage, we define the element’s area by creating a simple bounding box centered on the predicted point with the width and height equal to 14% of the sub-image’s width and height. We have also explored a more sophisticated approach using OmniParser to refine this bounding box. However, it does not improve performance and increases the inference overhead. In the selection stage, we use Qwen2.5VL-7B-instruct to choose the final answer and the prompt is listed in Fig. 4.
Default Hyperparameters.
Tab. 4 provides the default hyperparameters of UI-AGILE where cropping factor is the width and height ratio of new attempted image and last attempted image.
Hyperparameter Value
learning rate from 1e-06 to 4.36e-10
num generations 8
num train epochs 2
per device train batch size 4
cropping factor 0.6
gradient accumulation steps 4
Table 4: Default Hyperparameters
Appendix B Code
We provide the code for our RFT training and the Decomposed Grounding with Selection method in two separate modules. To avoid potential dependency conflicts, each module is designed to be run in its own conda environment.
To ensure a fair and comprehensive comparison, we conducted extensive experiments on the ScreenSpot-Pro benchmark. This involved re-implementing baseline models and evaluating them with our Decomposed Grounding with Selection method.
We use the parquet format to store test data in order to reduce the I/O read overhead.
We still have room for improvement in the implementation of Decomposed Grounding with Selection for Inference, including multi-threading to accelerate processing images and other non-GPU operations.
Figure 4: Prompt for VLM-based adjudication.
Appendix C Ablation Study
Fig 5 provides results of the ablation study using Type and GR as evaluation metrics on AndroidControl. We can observe that each component in UI-AGILE indeed contributes to the overall performance.
Figure 5: Ablation study using Type and GR as evaluation metrics on AndroidControl.
Appendix D Analysis of Attempts Per Step
Fig. 6 shows the distribution of attempts per GRPO training step, where each step processes a batch of two training samples. In the first epoch, we find that only 61.8% of training steps are fully successful on the initial attempt (i.e., both samples in the batch are solved without resampling). This means that without our strategy, a minimum of 19.1% ($38.2\%\div 2$) of training samples would have provided no learning signal. Overall attempt numbers decreases in the second epoch, demonstrating that the model learns from the samples salvaged by our method.